OpenStudy (anonymous):

-18/32 / 22/25 simplify and show work

OpenStudy (anonymous):

is the answer -45/77 show work

jhonyy9 (jhonyy9):

-18/32 = -9/16 so -9/16 / 22/25 = -9/16 * 25/22 = (-9*25)/(16*22) = - 225/352